# Equivariant and Invariant Parametrized Topological Complexity

### v3

@@
Abstract.
For a $G$-equivariant fibration $p\colon E\to B$, we introduce and study the invariant analogue of Cohen, Farber and Weinberger’s parametrized topological complexity, called the invariant parametrized topological complexity. This notion generalizes the invariant topological complexity introduced by Lubawski and Marzantowicz. We establish the equivariant fibrewise homotopy invariance of this notion and derive several bounds, including a cohomological lower bound and a dimensional upper bound. Additionally, we compare invariant parametrized topological complexity with other well-known invariants. When $G$ is a compact Lie group acting freely on $E$, we show that the invariant parametrized topological complexity of the $G$-fibration $p\colon E\to B$ coincides with the parametrized topological complexity of the induced fibration $\overline{p}\colon\overline{E}\to\overline{B}$ between the orbit spaces. Finally, we compute the invariant parametrized topological complexity of equivariant Fadell-Neuwirth fibrations, which measures the complexity of motion planning in presence of obstacles having unknown positions such that the order in which they are placed is irrelevant.
Apart from this, we establish several bounds, including a cohomological lower bound, an equivariant homotopy dimension-connectivity upper bound and various product inequalities for the equivariant sectional category. Applying them, we obtain some interesting results for equivariant and invariant parametrized topological complexity of a $G$-fibration.
Key words and phrases:
equivariant parametrized topological complexity, invariant topological complexity, equivariant sectional category, Lusternik-Schnirelmann category
2020 Mathematics Subject Classification:
55M30, 55R91, 55S40
@@ 1. Introduction
where $PX$ denotes the free path space of $X$ equipped with the compact-open topology. The concept of topological complexity was introduced by Farber in [16] to analyze the computational challenges associated with motion planning algorithms for the configuration space $X$ of a mechanical system. Over the past two decades, this invariant has attracted significant attention and has been a subject of extensive research.
@@
Recently, a novel parametrized approach to the theory of motion planning algorithms was introduced in [8, 9]. This approach provides enhanced universality and flexibility, allowing motion planning algorithms to operate effectively in diverse scenarios by incorporating external conditions. These external conditions are treated as parameters and form an integral part of the algorithm’s input. A parametrized motion planning algorithm takes as input a pair of configurations subject to the same external conditions and produces a continuous motion of the system that remains consistent with these external conditions.
We now briefly define the concept of parametrized topological complexity. For a fibration $p\colon E\to B$, let $E\times_{B}E$ denote the fibre product, which is the space of all pair of points in $E$ that lie in a common fibre of $p$. Let $E^{I}_{B}$ denote the space of all paths in $E$ whose images are contained within in a single fibre. Define the parametrized endpoint map
@@
In [8], it is shown that $\Pi$ is a fibration. The parametrized topological complexity of a fibration $p\colon E\to B$, denoted by $\mathrm{TC}[p\colon E\to B]$, is the smallest positive integer $k$ such that there is an open cover $\{U_{1},\dots,U_{k}\}$ of $E\times_{B}E$, where each $U_{i}$ admits a continuous section of $\Pi$. For further details and interesting computational results for parametrized topological complexity, see [8], [9], [18] and [29]. Additionally, the concept has been extended to fibrewise spaces by García-Calcines in [19]. On the other hand, Crabb [11] established some computational results in the fibrewise setting.
One of the key motivations for introducing this concept was to address the challenge of collision-free motion planning in environments where obstacles have unknown positions in advance. This can be described by the following scenario: A military commander oversees a fleet of $t$ submarines navigating waters with $s$ mines. The positions of these mines change every 24 hours. Each day, the commander must determine a movement plan for each submarine, ensuring that they travel from their current locations to their designated destinations without colliding with either the mines or other submarines. A parametrized motion planning algorithm will take as input the positions of the mines and the current and the desired positions of the submarines and will produced as output a collision-free motion of a fleet. Hence, the complexity of the universal motion planning algorithm in this setting can be described as the parametrized topological complexity of the Fadell-Neuwirth fibration
@@
where $F(\mathbb{R}^{d},s)$ is the configuration space of $s$ distinct points lying in $\mathbb{R}^{d}$, see Section 5.
@@
for any $\sigma$ in the permutation group $\Sigma_{s}$, the military commander should assign the same motion plan for the submarines. This is because both configurations describe the mines being placed at the same set of positions, regardless of their labeling. Thus, we should consider the unordered configuration space $F(\mathbb{R}^{d},s)/\Sigma_{s}$ for the placement of mines. Hence, in this new perspective, the complexity of the universal motion planning algorithm should be described as the parametrized topological complexity of the induced fibration
$\overline{p}\colon\overline{F(\mathbb{R}^{d},s+t)}\to\overline{F(\mathbb{R}^{d},s)}$
which is obtained by $p$ by taking the quotient under the natural action of $\Sigma_{s}$ on the configuration spaces. In this paper, we introduce the notion of invariant parametrized topological complexity for a $G$-fibration $p\colon E\to B$, denoted by $\mathrm{TC}^{G}[p\colon E\to B]$, to measure the complexity of parametrized motion planning problem where the order in which the mines are placed is irrelevant.
The invariant parametrized topological complexity is a parametrized analogue of the invariant topological complexity, which was introduced by Lubawski and Marzantowicz [25]. The invariant topological complexity for a $G$-space $X$, denoted by $\mathrm{TC}^{G}(X)$, behaves well with respect to quotients. In particular, if a compact Lie group $G$ acts freely on $X$, then the equality $\mathrm{TC}^{G}(X)=\mathrm{TC}(X/G)$ holds (see [25, Theorem 3.10]). Generalizing this to the parameterized setting we establish the following theorem.
@@
The aim of this paper is twofold. First, we examine various properties of the equivariant sectional category and equivariant parametrized topological complexity. Using these properties, we develop and analyze the new concept of invariant parametrized topological complexity, which we introduce in Section 4.
@@
In Section 2.1, we study the equivariant sectional category of a $G$-fibration, and establish multiple lower bounds in Theorem 2.2, Proposition 2.4 and Proposition 2.7. We also provide an equivariant homotopy dimension-connectivity upper bound in Theorem 2.12. Afterwards, we establish product inequalities in Proposition 2.14 and Corollary 2.15.
@@
In Section 2.2, we recall the notion of the equivariant LS category of a $G$-space. In this section, we establish a lower bound in terms of fixed point sets, and provide an equivariant homotopy dimension-connectivity upper bound, as stated in Proposition 2.20 and Theorem 2.21, respectively.
@@
Subsequently, Section 2.3 devoted to the equivariant and invariant topological complexity of a $G$-space, and we provide an equivariant homotopy dimension-connectivity upper bound for the former in Theorem 2.23.
@@
In Section 3, we explore various properties of the equivariant parametrized topological complexity of $G$-fibrations $p\colon E\to B$. Our main result Theorem 3.2, characterizes the elements of parametrized motion planning cover as the $G$-compressible subsets of the fibre product $E\times_{B}E$ into the diagonal $\Delta(E)$. Furthermore, we establish some lower bounds and the product inequalities in Proposition 3.4, Theorem 3.6 and Theorem 3.7, respectively.
@@
In Section 4, we introduce the notion of invariant parametrized topological complexity for $G$-fibrations. We establish the fibrewise $G$-homotopy invariance of this notion, and show that it generalizes both the parametrized and invariant topological complexity; see Theorem 4.3 and Proposition 4.4, respectively. For a $G$-fibration $p\colon E\to B$, in Theorem 4.7, we show that the elements of invariant parametrized motion planning cover can be characterized as the $(G\times G)$-compressible subsets of the fibre product $E\times_{B/G}E$ into the saturated diagonal $\mathbb{k}(E)=E\times_{E/G}E$. In Section 4.1, we investigate various properties and bounds for $\mathrm{TC}^{G}[p\colon E\to B]$. For example, we establish inequality under pullbacks (Proposition 4.9), dimensional upper bound (Proposition 4.10), lower bound (Proposition 4.13), cohomological lower bounds (Theorem 4.17 and Theorem 4.19), and product inequality (Theorem 4.20). Finally, we prove one of our main result, Theorem 4.29, which shows that the $\mathrm{TC}^{G}[p\colon E\to B]$ coincides with the parametrized topological complexity of the corresponding orbit fibration, when $G$ acts freely on $E$.
@@
In Section 5, we compute the invariant parametrized topological complexity of the equivariant Fadell-Neuwirth fibrations. Specifically, in Theorem 5.6 and Theorem 5.11, we establish:
@@
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=\begin{cases}2t+s,&\text{if $d$ is odd},\\ \text{either }2t+s-1\text{ or }2t+s&\text{if $d$ is even}.\end{cases}$
@@
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=2t+s-1.$
@@ 2.1. Equivariant sectional category
Schwarz [32] introduced and studied the notion of sectional category of a fibration, and later by Bernstein and Ganea in [3] for any map. The corresponding equivariant analogue was introduced by Colman and Grant in [10].
Definition 2.1 ([10, Definition 4.1]).
@@
First we establish a cohomological lower bound on the equivariant sectional category of a $G$-map using Borel cohomology. To the best of our knowledge, such a bound has not been documented in the literature. We believe that this result must already be known to experts in the field. Nevertheless, we provide a thorough proof of this result here.
@@
Suppose $\mathrm{secat}_{G}(p)\leq k$. Then there exists a $G$-invariant open cover $\{U_{1},\dots,U_{k}\}$ of $B$ such that each $U_{i}$ admits a $G$-equivariant homotopy section $s_{i}$ of $p$. Let $j_{i}\colon U_{i}\hookrightarrow B$ be the inclusion map. Then
$((j_{i})^{h}_{G})^{*}(u_{i})=((s_{i})^{h}_{G})^{*}\left((p^{h}_{G})^{*}(u_{i})\right)=0$
since $p\circ s_{i}\simeq_{G}j_{i}$ implies $((j_{i})^{h}_{G})^{*}=((s_{i})^{h}_{G})^{*}\circ(p^{h}_{G})^{*}$. Hence, the long exact sequence in cohomology associated to the pair $(B^{h}_{G},(U_{i})^{h}_{G})$ gives an element $v_{i}\in H^{*}(B^{h}_{G},(U_{i})^{h}_{G};R)$ such that $((q_{i})^{h}_{G})^{*}(v_{i})=u_{i}$, where $q_{i}\colon B\hookrightarrow(B,U_{i})$ is the inclusion map. Hence, we get
@@
Moreover, by the naturality of cup products, we have $(q^{h}_{G})^{*}(v_{1}\smile\cdots\smile v_{k})=u_{1}\smile\cdots\smile u_{k}$, where $q\colon B\hookrightarrow(B,B)$ is the inclusion map. Hence, $u_{1}\smile\cdots\smile u_{k}=0$. ∎
@@
  1. (1)
    Observe that if $G$ acts trivially on $X$, then the lower bound in Theorem 2.2 recovers the cohomological lower bound given by Schawrz in [32, Theorem 4].
  2. (2)
    Note the following commutative diagram of $G$-maps
    ${X}$${PX}$${X\times X,}$$\scriptstyle{\Delta}$$\scriptstyle{h}$$\scriptstyle{\pi}$
    where $h$ is a $G$-homotopy equivalence. Then, the lower bound in Theorem 2.2 recovers the bound [10, Theorem 5.15] on $\mathrm{TC}_{G}(X)$, obtained by Colman and Grant. More generally, it also recovers the bound [12, Theorem 4.25] on the equivariant parametrized topological complexity, obtained by the second author.
@@
[…]
@@
[…]
@@
Suppose $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration. Then $\overline{p}\colon\overline{E}\to\overline{B}$ is a fibration. Furthermore, if $E$ and $B$ are Hausdorff, and $G$ acts freely on $B$, then
@@
By [20, Corollary 2], it follows that $\overline{p}$ is a fibration. Note that the inequality $\mathrm{secat}(\overline{p})\leq\mathrm{secat}_{G}(p)$ follows from Proposition 2.4. Now we will show the reverse inequality. Note that $G$ also acts freely on $E$ since $p$ is a $G$-map. Hence, the diagram (3) is a pullback in the category of $G$-spaces, see the proof of [7, Theorem II.7.3]. Suppose $\overline{U}$ is an open subset of $\overline{B}$ and $\overline{s}\colon\overline{U}\to\overline{E}$ is a section of $\overline{p}$. Let $U=\pi_{B}^{-1}(\overline{U})$. Then, by the universal property of pullbacks, there exists a unique $G$-map $s\colon U\to E$ such that $\pi_{E}\circ s=\overline{s}\circ\pi_{B}\colon U\to\overline{E}$ and $p\circ s=i_{U}\colon U\to B$. Hence, $\mathrm{secat}_{G}(p)\leq\mathrm{secat}(\overline{p})$. ∎
@@
For any space $X$, the free path space fibration $\pi\colon PX\to X\times X$ is a $\mathbb{Z}_{2}$-fibration, with $\mathbb{Z}_{2}$-action on $PX$ given by reversal of paths and on $X\times X$ by transposition of factors, see [21, Example 2.6]. Hence, by Theorem 2.5, we get $\mathrm{secat}(\overline{\pi})\leq\mathrm{secat}_{\mathbb{Z}_{2}}(\pi)$. Thus, we recover the cohomological lower bound on the symmetrized topological complexity $\mathrm{TC}^{\Sigma}(X)$ in [21, Theorem 4.6] since the following commutative diagram
@@
Suppose $X$ is a $G$-space. For a subgroup $H$ of $G$, define the $H$-invariant subspace of $X$ as
@@
Suppose $p\colon E\to B$ is a $G$-fibration. If $H$ and $K$ are subgroups of $G$ such that the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a $K$-map, then
$\mathrm{secat}_{K}(p^{H})\leq\mathrm{secat}_{G}(p).$
Proof.
Suppose $s\colon U\to E$ is a $G$-equivariant section of $p$. Define $V=U\cap E^{H}$. Note that $V$ is $H$-fixed points of $U$ and is $K$-invariant. Since $s$ is $G$-equivariant, it takes $H$-fixed points to $H$-fixed points, and hence restricts to a $K$-equivariant map $\left.s\right|_{V}\colon V\to B^{H}$. It is clear that $\left.s\right|_{V}$ is a section of $p^{H}$. ∎
Corollary 2.8.

### v4

@@
Abstract
For a $G$-equivariant fibration $p\colon E\to B$, we introduce and study the invariant analogue of Cohen, Farber and Weinberger’s parametrized topological complexity, called the invariant parametrized topological complexity. This notion generalizes the invariant topological complexity introduced by Lubawski and Marzantowicz. When $G$ is a compact Lie group acting freely on $E$, we show that the invariant parametrized topological complexity of the $G$-fibration $p\colon E\to B$ coincides with the parametrized topological complexity of the induced fibration $\overline{p}\colon\overline{E}\to\overline{B}$ between the orbit spaces. Furthermore, we compute the invariant parametrized topological complexity of equivariant Fadell–Neuwirth fibrations, which measures the complexity of motion planning in the presence of obstacles with unknown positions, where the order of their placement is irrelevant. In addition, we study the equivariant sectional category and the equivariant parametrized topological complexity, which serve as essential tools for obtaining several results in this paper.
Keywords: equivariant parametrized topological complexity, invariant topological complexity, equivariant sectional category, Lusternik-Schnirelmann category
2020 Mathematics Subject Classification: 55M30, 55R91, 55S40
@@ 1. Introduction
where $PX$ denotes the free path space of $X$ equipped with the compact-open topology. The concept of topological complexity was introduced by Farber in [FarberTC] to analyze the computational challenges associated with motion planning algorithms for the configuration space $X$ of a mechanical system. Over the past two decades, this invariant has attracted significant attention and has been a subject of extensive research.
@@
Recently, a novel parametrized approach to the theory of motion planning algorithms was introduced in [farber-para-tc, PTCcolfree]. This approach provides enhanced universality and flexibility, allowing motion planning algorithms to operate effectively in diverse scenarios by incorporating external conditions. These external conditions are treated as parameters and form an integral part of the algorithm’s input. A parametrized motion planning algorithm takes as input a pair of configurations subject to the same external conditions and produces a continuous motion of the system that remains consistent with these external conditions.
We now briefly define the concept of parametrized topological complexity. For a fibration $p\colon E\to B$, let $E\times_{B}E$ denote the fibre product, which is the space of all pair of points in $E$ that lie in a common fibre of $p$. Let $E^{I}_{B}$ denote the space of all paths in $E$ whose images are contained within a single fibre. Define the parametrized endpoint map
@@
In [farber-para-tc], it is shown that $\Pi$ is a fibration. The parametrized topological complexity of a fibration $p\colon E\to B$, denoted by $\mathrm{TC}[p\colon E\to B]$, is the smallest positive integer $k$ such that there is an open cover $\{U_{1},\dots,U_{k}\}$ of $E\times_{B}E$, where each $U_{i}$ admits a continuous section of $\Pi$. For further details and interesting computational results for parametrized topological complexity, see [farber-para-tc], [PTCcolfree], [ptcspherebundles] and [minowa2024parametrized]. Additionally, the concept has been extended to fibrewise spaces by García-Calcines in [fibrewise]. On the other hand, Crabb [crabb2023fibrewise] established some computational results in the fibrewise setting.
One of the key motivations for introducing this concept was to address the challenge of collision-free motion planning in environments where obstacles have unknown positions in advance. This can be described by the following scenario, as illustrated in [farber-para-tc]: A military commander oversees a fleet of $t$ submarines navigating waters with $s$ mines. The positions of these mines change every 24 hours. Each day, the commander must determine a motion plan for each submarine, ensuring that they travel from their current locations to their designated destinations without colliding with either the mines or other submarines. A parametrized motion planning algorithm will take as input the positions of the mines on the given day and the current and the desired positions of the submarines and will produce as output a collision-free motion of a fleet. Hence, the complexity of the universal motion planning algorithm in this setting can be described as the parametrized topological complexity of the Fadell-Neuwirth fibration
@@
where $F(\mathbb{R}^{d},s)$ is the configuration space of $s$ distinct points lying in $\mathbb{R}^{d}$, see Section˜5.
@@
for any $\sigma$ in the permutation group $\Sigma_{s}$, the military commander should assign the same motion plan for the submarines. This is because both configurations describe the mines being placed at the same set of positions, regardless of their labeling. Thus, we should consider the unordered configuration space $F(\mathbb{R}^{d},s)/\Sigma_{s}$ for the placement of mines. Hence, in this new perspective, the complexity of the universal motion planning algorithm should be described as the parametrized topological complexity of the fibration
$\overline{p}\colon\overline{F(\mathbb{R}^{d},s+t)}\to\overline{F(\mathbb{R}^{d},s)},$
which is induced from $p$ by taking the quotient under the natural action of $\Sigma_{s}$ on the configuration spaces.
In this paper, we introduce the notion of invariant parametrized topological complexity for a $G$-fibration $p\colon E\to B$, denoted by $\mathrm{TC}^{G}[p\colon E\to B]$, which seems to provide a more suitable framework for analyzing the impact of symmetries on parametrized motion planning algorithms. The invariant parametrized topological complexity is a parametrized analogue of the invariant topological complexity, which was introduced by Lubawski and Marzantowicz [invarianttc]. The invariant topological complexity for a $G$-space $X$, denoted by $\mathrm{TC}^{G}(X)$, behaves well with respect to quotients. In particular, if a compact Lie group $G$ acts freely on $X$, then the equality $\mathrm{TC}^{G}(X)=\mathrm{TC}(X/G)$ holds (see [invarianttc, Theorem 3.10]). Generalizing this to the parameterized setting we establish the following theorem.
@@
The aim of this paper is twofold. First, we examine various properties of the equivariant sectional category and equivariant parametrized topological complexity. Using these properties, we develop and analyze the new concept of invariant parametrized topological complexity, which we introduce in Section˜4.
In Section˜2.1, we introduce the concept of equivariant fibrations and present several examples, including equivariant covering maps.
@@
In Section˜2.2, we study the equivariant sectional category of a $G$-fibration, and establish multiple lower bounds in Theorem˜2.7, Proposition˜2.9 and Proposition˜2.12. We also provide an equivariant homotopy dimension-connectivity upper bound in Theorem˜2.16. Afterwards, we establish product inequalities in Proposition˜2.18 and Corollary˜2.19.
@@
In Section˜2.3, we recall the notion of the equivariant LS category of a $G$-space. In this section, we establish a lower bound in terms of fixed point sets, and provide an equivariant homotopy dimension-connectivity upper bound, as stated in Proposition˜2.24 and Theorem˜2.26, respectively.
@@
Subsequently, Section˜2.4 devoted to the equivariant and invariant topological complexity of a $G$-space, and we provide an equivariant homotopy dimension-connectivity upper bound for the former in Theorem˜2.28.
@@
In Section˜3, we explore various properties of the equivariant parametrized topological complexity of $G$-fibrations $p\colon E\to B$. Our main result Theorem˜3.2, characterizes the elements of a parametrized motion planning cover as the $G$-compressible subsets of the fibre product $E\times_{B}E$ into the diagonal $\Delta(E)$. Furthermore, we establish some lower bounds and the product inequalities in Proposition˜3.4, Theorem˜3.6 and Theorem˜3.8, respectively.
@@
In Section˜4, we introduce the notion of invariant parametrized topological complexity for $G$-fibrations. We establish the fibrewise $G$-homotopy invariance of this notion, and show that it generalizes both the parametrized and invariant topological complexity; see Theorem˜4.3 and Proposition˜4.4, respectively. For a $G$-fibration $p\colon E\to B$, in Theorem˜4.6, we show that the elements of an invariant parametrized motion planning cover can be characterized as the $(G\times G)$-compressible subsets of the fibre product $E\times_{B/G}E$ into the saturated diagonal $\mathbb{k}(E)=E\times_{E/G}E$. In Section˜4.1, we investigate various properties and bounds for $\mathrm{TC}^{G}[p\colon E\to B]$. For example, we establish inequality under pullbacks (Proposition˜4.8), dimensional upper bound (Proposition˜4.9), lower bound (Proposition˜4.12), cohomological lower bounds (Theorem˜4.15 and Theorem˜4.16), and product inequality (Theorem˜4.17). Finally, we prove one of our main results, Theorem˜4.26, which shows that the $\mathrm{TC}^{G}[p\colon E\to B]$ coincides with the parametrized topological complexity of the corresponding orbit fibration, when $G$ acts freely on $E$. We conclude the section with some examples illustrating the application of Theorem˜4.26.
@@
In Section˜5, we compute the invariant parametrized topological complexity of the equivariant Fadell-Neuwirth fibrations. Specifically, in Theorem˜5.6 and Theorem˜5.10, we establish:
@@
[…]
@@
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{2},s+t)\to F(\mathbb{R}^{2},s)]=2t+s-1.$
Notations and conventions
Throughout the text, $G$ denotes a general topological group acting on Hausdorff spaces, unless stated otherwise. We adopt standard terminology and notation from equivariant topology, such as $G$-spaces, $G$-maps, $G$-homotopies and related notions. In this paper, all category-type invariants are taken to be un-normalized, that is, they equal the number of open sets in the cover. Thus, our definitions exceed by one those in [farber-para-tc, PTCcolfree], but agree with those used in [FarberTC, colmangranteqtc, invarianttc].
@@ 2. Preliminaries
2.1. $G$-Fibrations
We begin by recalling the definition of a $G$-fibration, followed by a few examples. For a more detailed discussion, we refer the reader to [gevorgyan2023equivariant] and [grant2019symmetrized, Section 2].
Definition 2.1.
A $G$-map $p\colon E\to B$ is called a $G$-fibration if it has the $G$-homotopy lifting property with respect to any $G$-space $X$. More precisely, if $H\colon X\times I\to B$ is a $G$-homotopy and $h\colon X\to E$ is a $G$-map with $p\circ h=H_{0}$, then there exists a $G$-homotopy $\widetilde{H}\colon X\times I\to E$ that satisfies $p\circ\widetilde{H}=H$ and $\widetilde{H}_{0}=h$.
Example 2.2.
Here we list some examples of $G$-fibrations.
  1. (1)
    For any $G$-spaces $B$ and $F$, the projection maps $\pi_{1}\colon B\times F\rightarrow B$ and $\pi_{2}\colon B\times F\to F$ are $G$-fibrations.
  2. (2)
    If $G$ is a compact Hausdorff topological group, then every principal $G$-bundle over a paracompact space is a $G$-fibration, where $G$ acts trivially on the base. This can be shown using [gevorgyan2023equivariant, Theorem 7]. In particular, the Hopf fibration $S^{1}\hookrightarrow S^{3}\to S^{2}$ is a $S^{1}$-fibration.
  3. (3)
    For any space $X$, the free path space fibration $\pi\colon PX\to X\times X$ is a $\mathbb{Z}_{2}$-fibration, with $\mathbb{Z}_{2}$-action on $PX$ given by reversal of paths and on $X\times X$ by transposition of factors, see [grant2019symmetrized, Example 2.6].
  4. (4)
    For any $G$-space $X$, the free path space fibration is a $G$-fibration. Moreover, if $x_{0}\in X$ is a fixed point under $G$-action, then the path space fibration is also a $G$-fibration. See Section˜2.4 and Section˜2.3 respectively for more details.
The following theorem provides a sufficient condition for a fibration to be a $G$-fibration. We note that although the proof is presented for compact Hausdorff groups, it remains valid for arbitrary topological groups.
Theorem 2.3 ([gevorgyan2023equivariant, Lemma 5]).
If $p\colon E\to B$ is a $G$-map that is a fibration with the unique path-lifting property, then $p$ is a $G$-fibration.
The following is an immediate corollary of this theorem.
Corollary 2.4.
If $p\colon E\to B$ is a $G$-map that is a covering map, then $p$ is a $G$-fibration. In particular, if $p\colon\widetilde{X}\to X$ is a universal covering, then $p$ is a $\pi_{1}(X)$-fibration, where $\pi_{1}(X)$ acts on $\widetilde{X}$ via deck transformations and acts trivially on $X$.
Example 2.5.
Here we list some non-trivial examples of $G$-fibrations.
  1. (1)
    Suppose $\alpha\in\mathbb{R}$. Then the universal covering $p\colon\mathbb{R}\to S^{1}$, given by $t\mapsto e^{2\pi it}$, is a $\mathbb{Z}$-fibration, where $\mathbb{Z}$ acts on $\mathbb{R}$ by $n\cdot t=t+n\alpha$, and acts on $S^{1}$ by $n\cdot z=e^{2\pi in\alpha}z$.
  2. (2)
    The universal covering $p\colon S^{n}\to\mathbb{R}P^{n}$ is a $\Sigma_{n+1}$-fibration, where the symmetric group $\Sigma_{n+1}$ acts on $S^{n}$ by $\sigma\cdot(x_{0},\dots,x_{n})=(x_{\sigma(0)},\dots,x_{\sigma(n)}),$ and acts on $\mathbb{R}P^{n}$ by $\sigma\cdot[x_{0}:\dots:x_{n}]=[x_{\sigma(0)}:\dots:x_{\sigma(n)}].$
@@ 2.2. Equivariant sectional category
Schwarz [Sva] introduced and studied the notion of sectional category of a fibration, and later by Berstein and Ganea in [secat] for any map. The corresponding equivariant analogue was introduced by Colman and Grant in [colmangranteqtc].
Definition 2.6 ([colmangranteqtc, Definition 4.1]).
@@
First we establish a cohomological lower bound on the equivariant sectional category of a $G$-map using Borel cohomology. This is based on Colman and Grant’s work [colmangranteqtc, Theorem 5.15], which provides a similar cohomological lower bound on equivariant topological complexity. To the best of our knowledge, such a bound has not been documented in the literature. We believe that this result must already be known to experts in the field. Nevertheless, we provide a thorough proof of this result here.
@@
Suppose $\mathrm{secat}_{G}(p)\leq k$. Then there exists a $G$-invariant open cover $\{U_{1},\dots,U_{k}\}$ of $B$ such that each $U_{i}$ admits a $G$-equivariant homotopy section $s_{i}$ of $p$. If $j_{i}\colon U_{i}\hookrightarrow B$ is the inclusion map, then $((j_{i})^{h}_{G})^{*}(u_{i})=((s_{i})^{h}_{G})^{*}\left((p^{h}_{G})^{*}(u_{i})\right)=0.$ Hence, the long exact sequence in cohomology of the pair $(B^{h}_{G},(U_{i})^{h}_{G})$ yields an element $v_{i}\in H^{*}(B^{h}_{G},(U_{i})^{h}_{G};R)$ such that $((q_{i})^{h}_{G})^{*}(v_{i})=u_{i}$, where $q_{i}\colon B\hookrightarrow(B,U_{i})$ is the inclusion map. Consequently, we obtain
@@
Therefore, by the naturality of the cup product, we have $u_{1}\smile\cdots\smile u_{k}=(q^{h}_{G})^{*}(v_{1}\smile\cdots\smile v_{k})=0$, where $q\colon B\hookrightarrow(B,B)$ is the inclusion map. ∎
@@
Observe that if $G$ acts trivially on $X$, then the lower bound in Theorem˜2.7 coincides with the cohomological lower bound given by Schwarz in [Sva, Theorem 4]. Moreover, it generalizes the lower bounds for equivariant topological complexity and equivariant parametrized topological complexity previously established by Colman–Grant [colmangranteqtc, Theorem 5.15] and Daundkar [D-EqPTC, Theorem 4.25], respectively.
@@
$E¯{\overline{E}}B¯,{\overline{B},}$$\scriptstyle{\pi_{E}}$$\scriptstyle{p}$$\scriptstyle{\pi_{B}}$$\scriptstyle{\overline{p}}$ (3)
@@
${\overline{E}}U¯{\overline{U}}$$\scriptstyle{\pi_{E}\circ s}$$\scriptstyle{\pi_{B}}$$\scriptstyle{\overline{s}}$
@@
Suppose $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration. Then $\overline{p}\colon\overline{E}\to\overline{B}$ is a fibration. Furthermore, if $G$ acts freely on $B$, then
@@
By [gevorgyan2023equivariant, Corollary 2], it follows that $\overline{p}$ is a fibration. Note that the inequality $\mathrm{secat}(\overline{p})\leq\mathrm{secat}_{G}(p)$ follows from Proposition˜2.9. Now we will show the reverse inequality. Note that $G$ acts freely on $E$ as well, since $G$ acts freely on $B$ and $p$ is a $G$-map. If $\phi\colon E\to\overline{E}\times_{\overline{B}}B$ denotes the natural $G$-map induced by the universal property of the pullback, then $\phi$ preserves the orbit structure because $G$ acts freely on both $B$ and $E$. Hence, the diagram (3) is a pullback in the category of $G$-spaces, see the proof of [bredon-transformation-groups, Theorem II.7.3]. Suppose $\overline{U}$ is an open subset of $\overline{B}$ and $\overline{s}\colon\overline{U}\to\overline{E}$ is a section of $\overline{p}$. Let $U=\pi_{B}^{-1}(\overline{U})$. Then, by the universal property of pullbacks, there exists a unique $G$-map $s\colon U\to E$ such that $\pi_{E}\circ s=\overline{s}\circ\pi_{B}\colon U\to\overline{E}$ and $p\circ s=i_{U}\colon U\to B$. Hence, $\mathrm{secat}_{G}(p)\leq\mathrm{secat}(\overline{p})$. ∎
@@
As mentioned in Example˜2.2 (3), the free path space fibration $\pi\colon PX\to X\times X$ is a $\mathbb{Z}_{2}$-fibration for any space $X$. The $\mathbb{Z}_{2}$-action on $PX$ is given by reversal of paths, and on $X\times X$ it is given by transposition of factors. Hence, by Theorem˜2.10, we get $\mathrm{secat}(\overline{\pi})\leq\mathrm{secat}_{\mathbb{Z}_{2}}(\pi)$. Thus, we recover the cohomological lower bound on the symmetrized topological complexity $\mathrm{TC}^{\Sigma}(X)$ in [grant2019symmetrized, Theorem 4.6], since the following commutative diagram
@@
Suppose $X$ is a $G$-space. For a subgroup $H$ of $G$, define the $H$-fixed subspace of $X$ as
@@
Suppose $p\colon E\to B$ is a $G$-fibration. If $H$ and $K$ are subgroups of $G$ such that $E^{H}$ and $B^{H}$ are $K$-invariant, and the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a $K$-map, then
$\mathrm{secat}_{K}(p^{H})\leq\mathrm{secat}_{G}(p).$ (4)
In particular, if $G$ is a compact Hausdorff topological group, then